# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v1

@@ 1. Introduction
Quantum error correction plays a central role in the development of quantum computing architectures. With the rapid advancement of technologies involving quantum computation and communication, there is a growing demand for effective quantum error-correcting codes (QECCs). These codes are essential for protecting quantum information from noise, which can arise from various types of quantum channels. In the celebrated work of Calderbank, Shor, and Steane [6, 38], the authors proposed the CSS construction that utilizes the duality of certain classical error-correcting codes to construct quantum error-correcting codes. This method is commonly used to construct QECCs over finite fields [23]. QECCs from the CSS construction are most often constructed using self-orthogonal error-correcting codes and its dual, where self-orthogonal codes have applications in lattice theory [42], row-self-orthogonal matrices [29, 30], linear complementary dual codes [30], and many others. Beyond self-orthogonality, the notions of dual-containing and self-dual codes are also significant. Self-dual codes notably exhibit deep connections with group theory [18, 22], lattices [35], and design theory [36], and are recognized for being asymptotically good [28]. Because of these applications, it is a worthwhile endeavor to determine the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes.
@@
The focus of this paper is on constructing quasi-cyclic (QC) error-correcting codes that are self-orthogonal and self-dual under the Euclidean and Hermitian inner products. Quasi-cyclic codes are a natural generalization of cyclic codes and are well-studied in classical coding theory literature. Quasi-cyclic codes have a rich algebraic structure [9, 16, 26] and have strong connections with convolutional codes [37]. Moreover, there exist QC codes that are asymptotically good [43]. In 2022, the authors in [31] found an upper bound for the locality of QC locally recoverable codes using generator matrices in systematic form where the first block is the identity matrix and the remaining blocks are comprised of smaller circulant matrices. More recently in  [39], a more concise way of finding an upper bound on the locality of these codes was studied, and a construction of optimal locally recoverable QC codes was presented.
@@
Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in [1, 12, 33], the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. Consequently, as these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
@@
In this paper, we propose a construction of infinite families of self-orthogonal and self-dual QC codes using their constituent code representation described in [16]. We then describe how the codes in our construction yield the existence of QECCs with good parameters. The paper is organized as follows: Section 2 provides background results related to QC codes and their constituent codes. This section also includes the foundations of constructing QECCs from classical ones. In Section 3, we present the main results of this paper, namely, the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (or self-dual) with respect to the Euclidean inner product. We show that the self-orthogonal QC codes in this construction satisfy the square-root-like lower bound to their minimum distance. We also introduce the notion of Galois-closed codes to obtain QC codes that are self-orthogonal with respect to the Hermitian inner product. Lastly in Section 4, we show how some codes from our construction are used to obtain quantum error-correcting codes with the best known parameters so far. All computations throughout this paper have been done using SageMath or Mathematica.
@@ 2.1. Quasi-Cyclic Codes and Their Representations
and the code is called a maximum distance separable code, or an MDS code, if equality in (1) is achieved. A quintessential example of a maximum distance separable code is the Reed-Solomon code described in [32]. More generally, if we choose nonzero elements $v_{1},...,v_{n}\in\mathbb{F}_{q}$ and distinct elements $\alpha_{1},...,\alpha_{n}\in\mathbb{F}_{q}$, we define the Generalized Reed-Solomon Codes as
@@
Moreover, as in [26], we consider the ring $R_{q,m}^{\ell}:=\left(\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle\right)^{\ell}$ where $\operatorname{gcd}(m,q)=1$. To any codeword c as in (3), we associate an element of $R_{q,m}^{\ell}$:
@@
By Lemma 3.1 in [26], the map $\phi:\mathbb{F}_{q}^{m\ell}\to R_{q,m}^{\ell}$ defined by $\phi(\textbf{c})=\vec{\textbf{c}}(x)$ is an $R$-module isomorphism. Hence, any QC code over $\mathbb{F}_{q}$ of length $m\ell$ and index $\ell$ can be viewed as an $R$-submodule of $R_{q,m}^{\ell}$. Throughout this paper, we will move freely between the representations in (4) and (5).
Given a polynomial $f(x)=f_{n}x^{n}+f_{n-1}x^{n-1}+\cdots+f_{0}\in\mathbb{F}_{q}[x]$, its reciprocal polynomial, $f^{*}$, is given by $f^{*}(x)=x^{n}f(x^{-1})$. As seen in [26], we may decompose $x^{m}-1\in\mathbb{F}_{q}[x]$ into irreducible factors
@@
The following lemma gives an easy way to compute the dimension of a QC code $\mathcal{C}$ from its constituent codes. We slightly modified the notation of the following result from [39] to align with the notation in this paper.
@@
[39, Lemma 2] Let $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ be an $m\ell,k]$ QC code and let $C_{j}\subseteq\mathbb{F}_{q}^{\ell}(\alpha^{u_{j}})\cong\left(\frac{\mathbb{F}_{q}[x]}{\langle b_{j}\rangle}\right)^{\ell}$ be its respective constituent $\ell,k_{j}]$ codes. Suppose that the nontrivial constituent codes of $\mathcal{C}$ are given by $C_{1},\ldots,C_{h}$. Then,
@@
Let $s\in\mathbb{Z}_{m}$ and let $C_{s}=\{s,sq,sq^{2},...,sq^{r-1}\}\subset\mathbb{Z}_{m}$ be the $q$-cyclotomic coset of $s$, where $r$ is the smallest positive integer so that $sq^{r}\equiv s\mod m$. It is important to note that the $q$-cyclotomic cosets split $\mathbb{Z}_{m}.$ This classic concept is seen in constructions of cyclic codes, as well as in the next theorem (formally known as the trace representation [26] of QC codes) where we slightly modified the notation in the statement to better align with the notation in this paper. Moreover, if $\mathbb{F}_{q}$ is a field and $\mathbb{F}_{q^{s}}$ a a field extension of $\mathbb{F}_{q}$, the trace map $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}:\mathbb{F}_{q^{s}}\to\mathbb{F}_{q}$ is defined as
@@
Theorem 2.2.
[26, Theorem 5.1] Let $\mathbb{F}_{q}$ be a finite field and $\gcd(m,q)=1$. Then, for any $\ell$, the quasi-cyclic codes over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ are precisely given by the following construction. Decompose $x^{m}-1$ as per (7) . Write $\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}=F_{i}$, $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}=G_{j}^{\prime}$, and $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}=G_{j}^{\prime\prime}$. Let $U_{i}$ (resp. $V_{j}$ and $W_{j}$) denote the cyclotomic coset of $\mathbb{Z}/m\mathbb{Z}$ corresponding to $F_{i}$ (resp. $G_{j}^{\prime}$ and $G_{j}^{\prime\prime}$) and fix $u_{i}\in U_{i}$, $v_{j}\in V_{j}$, and $w_{j}\in W_{j}$. For each $i$, let $C_{i}$ be a code of length $\ell$ over $F_{i}$, and for each $j$, let $C_{j}^{\prime}$ be a code of length $\ell$ over $G_{j}^{\prime}$ and let $C_{j}^{\prime\prime}$ be a code of length $\ell$ over $G_{j}^{\prime\prime}$. For $\mathbf{x_{i}}\in C_{i}$, $\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime}$, and $\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}$, and for each $0\leq g\leq m-1$, let
@@
2.2. The Duals of Quasi-Cyclic Codes
Recently, QC codes have been used in constructions of QECCs where orthogonality conditions on QC codes play a central role [1, 12, 33] . To study the duality of codes in $\mathbb{F}_{q}^{m\ell}$, we define the following inner products: Let $\textbf{a},\textbf{b}\in\mathbb{F}_{q}^{m\ell}$ where
@@ 2.2. The Duals of Quasi-Cyclic Codes
[26, Proposition 3.2] Let $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m\ell}$. Then $(T^{k\ell}(\mathbf{a}))\cdot\mathbf{b}=0$ for all $0\leq k\leq m-1$ if and only if $\langle\phi(\mathbf{a}),\phi(\mathbf{b})\rangle_{H}=0$, where $\phi$ is the aforementioned $R$-module isomorphism.
@@
[26, Corollary 3.3] Let $\mathcal{C}$ be a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ and let $\phi(\mathcal{C})$ be its image in $R_{q,m}^{\ell}$ under $\phi$. Then $\phi(\mathcal{C})^{\perp}=\phi(\mathcal{C}^{\perp})$, where the dual in $\mathbb{F}_{q}^{m\ell}$ is taken with respect to the Euclidean inner product, while the dual in $R_{q,m}^{\ell}$ is taken with respect to the Hermitian inner product. In particular, a quasi-cyclic code $\mathcal{C}$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if only if $\phi(\mathcal{C})$ is self-dual over $R_{q,m}$ with respect to the Hermitian inner product.
@@
[26, Theorem 4.2] An $\ell$ quasi-cyclic code $\mathcal{C}$ of length $m\ell$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if and only if
@@
    As observed in [19, page 136], the irreducible factors $f_{i}$ which are self-reciprocal imply that the cardinality $q_{i}$ of each field extension $F_{i}$ is an even power of $q$, for all $i$. However, there are two exceptions. The first one, for all $m$ and $q$, is the field extension coming from the irreducible factor x1x-1 of $x^{m}-1$. The other exception occurs when $q$ is odd and $m$ is even in which $x+1$ is another self-reciprocal factor of $x^{m}-1$. In both exceptions, $q_{i}=q$ and we equip these fields with the Euclidean inner product. Otherwise, we equip $F_{i}$ with the Hermitian inner product.
@@
The next result extends [26, Theorem 4.2] to the dual of a QC code.
@@
[19, Proposition 7.3.5] Let $\mathcal{C}$ be a QC code with CRT decomposition as (10). Then its dual code $\mathcal{C}^{\perp}$ (under the Euclidean inner product) is of the form
@@
[17, Lemma 2.10] Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k,d_{2}]$ codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$, respectively, then $G_{1}|G_{2}]$ generates an $n_{1}+n_{2},k,d_{1}+d_{2}]$ code over $\mathbb{F}_{q}$ .
@@
In the special case when $C_{1}=C_{2}$, the code with generator matrix $G_{1}|G_{1}]$ will be denoted as $2C_{1}$, and such an argument may be naturally extended to $mC_{1}$, for any positive integer $m$. In this work, for any code $C$, we will refer to the code $mC$ as $m$-copies of $C$ which will play an important role in the next section.

### v2

@@ 1. Introduction
Quantum error correction plays a central role in the development of quantum computing architectures. With the rapid advancement of technologies involving quantum computation and communication, there is a growing demand for effective quantum error-correcting codes (QECCs). These codes are essential for protecting quantum information from noise, which can arise from various types of quantum channels. In the celebrated work of Calderbank, Shor, and Steane [Calderbank-Shor1996, Steane1996], the authors proposed the CSS construction that utilizes the duality of certain classical error-correcting codes to construct quantum error-correcting codes. This method is commonly used to construct QECCs over finite fields [Ketkar-Klappenecker-Kumar-Sarvepalli2006]. QECCs from the CSS construction are most often constructed using self-orthogonal error-correcting codes and its dual, where self-orthogonal codes have applications in lattice theory [W98], row-self-orthogonal matrices [Massey92, Massey98], linear complementary dual codes [Massey98], and many others. Beyond self-orthogonality, the notions of dual-containing and self-dual codes are also significant. Self-dual codes notably exhibit deep connections with group theory [Hoehn00, JLLX13], lattices [SloaneChapter99], and design theory [Sole90], and are recognized for being asymptotically good [MACWILLIAMS1972153]. Because of these applications, it is a worthwhile endeavor to determine the existence and constructions of self-orthogonal, dual-containing, and self-dual classical error-correcting codes.
@@
The focus of this paper is on constructing quasi-cyclic (QC) error-correcting codes that are self-orthogonal and self-dual under the Euclidean and Hermitian inner products. Quasi-cyclic codes are a natural generalization of cyclic codes and are well-studied in classical coding theory literature. Quasi-cyclic codes have a rich algebraic structure [seguin, cem, lingsole] and have strong connections with convolutional codes [SolomonTilborg]. Moreover, there exist QC codes that are asymptotically good [longcodes]. In 2022, the authors in [RB] found an upper bound for the locality of QC locally recoverable codes using generator matrices in systematic form where the first block is the identity matrix and the remaining blocks are comprised of smaller circulant matrices. More recently in  [BastosAlvarezFloresSalerno], a more concise way of finding an upper bound on the locality of these codes was studied, and a construction of optimal locally recoverable QC codes was presented.
@@
Very recently, some constructions of self-dual, dual-containing, and self-orthogonal QC codes have been explored to develop effective quantum error-correcting codes. For example in [Abdukhalikov-Bag-Panario2023, GalindoHernando, benjwal], the authors obtained quantum codes with good parameters by studying the generators of the dual of a QC code to determine the conditions where such codes are dual-containing, self-orthogonal, or self-dual. Consequently, as these properties are intrinsically related to the CSS construction, QECCs with good parameters are obtained.
@@
In this paper, we propose a construction of infinite families of self-orthogonal and self-dual QC codes using their constituent code representation described in [cem]. We then describe how the codes in our construction yield the existence of QECCs with good parameters. The paper is organized as follows: Section 2 provides background results related to QC codes and their constituent codes. This section also includes the foundations of constructing QECCs from classical ones. In Section 3, we present the main results of this paper, namely, the use of constituent codes of QC codes to construct infinite families of QC codes that are self-orthogonal (or self-dual) with respect to the Euclidean inner product. We show that the self-orthogonal QC codes in this construction satisfy the square-root-like lower bound to their minimum distance. We also introduce the notion of Galois-closed codes to obtain QC codes that are self-orthogonal with respect to the Hermitian inner product. Lastly in Section 4, we show how some codes from our construction are used to obtain quantum error-correcting codes with the best known parameters so far. All computations throughout this paper have been done using SageMath or Mathematica.
@@ 2.1. Quasi-Cyclic Codes and Their Representations
and the code is called a maximum distance separable code, or an MDS code, if equality in (1) is achieved. A quintessential example of a maximum distance separable code is the Reed-Solomon code described in [RS1958]. More generally, if we choose nonzero elements $v_{1},...,v_{n}\in\mathbb{F}_{q}$ and distinct elements $\alpha_{1},...,\alpha_{n}\in\mathbb{F}_{q}$, we define the Generalized Reed-Solomon Codes as
@@
Moreover, as in [lingsole], we consider the ring $R_{q,m}^{\ell}:=\left(\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle\right)^{\ell}$ where $\operatorname{gcd}(m,q)=1$. To any codeword c as in (3), we associate an element of $R_{q,m}^{\ell}$:
@@
By Lemma 3.1 in [lingsole], the map $\phi:\mathbb{F}_{q}^{m\ell}\to R_{q,m}^{\ell}$ defined by $\phi(\textbf{c})=\vec{\textbf{c}}(x)$ is an $R$-module isomorphism. Hence, any QC code over $\mathbb{F}_{q}$ of length $m\ell$ and index $\ell$ can be viewed as an $R$-submodule of $R_{q,m}^{\ell}$. Throughout this paper, we will move freely between the representations in (4) and (5).
Given a polynomial $f(x)=f_{n}x^{n}+f_{n-1}x^{n-1}+\cdots+f_{0}\in\mathbb{F}_{q}[x]$, its reciprocal polynomial, $f^{*}$, is given by $f^{*}(x)=x^{n}f(x^{-1})$. As seen in [lingsole], we may decompose $x^{m}-1\in\mathbb{F}_{q}[x]$ into irreducible factors
@@
The following lemma gives an easy way to compute the dimension of a QC code $\mathcal{C}$ from its constituent codes. We slightly modified the notation of the following result from [BastosAlvarezFloresSalerno] to align with the notation in this paper.
@@
[BastosAlvarezFloresSalerno, Lemma 2] Let $\mathcal{C}\subset\mathbb{F}_{q}^{m\ell}$ be an $m\ell,k]$ QC code and let $C_{j}\subseteq\mathbb{F}_{q}^{\ell}(\alpha^{u_{j}})\cong\left(\frac{\mathbb{F}_{q}[x]}{\langle b_{j}\rangle}\right)^{\ell}$ be its respective constituent $\ell,k_{j}]$ codes. Suppose that the nontrivial constituent codes of $\mathcal{C}$ are given by $C_{1},\ldots,C_{h}$. Then,
@@
Let $s\in\mathbb{Z}_{m}$ and let $C_{s}=\{s,sq,sq^{2},...,sq^{r-1}\}\subset\mathbb{Z}_{m}$ be the $q$-cyclotomic coset of $s$, where $r$ is the smallest positive integer so that $sq^{r}\equiv s\mod m$. It is important to note that the $q$-cyclotomic cosets split $\mathbb{Z}_{m}.$ This classic concept is seen in constructions of cyclic codes, as well as in the next theorem (formally known as the trace representation [lingsole] of QC codes) where we slightly modified the notation in the statement to better align with the notation in this paper. Moreover, if $\mathbb{F}_{q}$ is a field and $\mathbb{F}_{q^{s}}$ a a field extension of $\mathbb{F}_{q}$, the trace map $\operatorname{Tr}_{\mathbb{F}_{q^{s}}/\mathbb{F}_{q}}:\mathbb{F}_{q^{s}}\to\mathbb{F}_{q}$ is defined as
@@
Theorem $2.$
[lingsole, Theorem 5.1] Let $\mathbb{F}_{q}$ be a finite field and $\gcd(m,q)=1$. Then, for any $\ell$, the quasi-cyclic codes over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ are precisely given by the following construction. Decompose $x^{m}-1$ as per (7) . Write $\frac{\mathbb{F}_{q}[x]}{\langle f_{i}\rangle}=F_{i}$, $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}\rangle}=G_{j}^{\prime}$, and $\frac{\mathbb{F}_{q}[x]}{\langle g_{j}^{*}\rangle}=G_{j}^{\prime\prime}$. Let $U_{i}$ (resp. $V_{j}$ and $W_{j}$) denote the cyclotomic coset of $\mathbb{Z}/m\mathbb{Z}$ corresponding to $F_{i}$ (resp. $G_{j}^{\prime}$ and $G_{j}^{\prime\prime}$) and fix $u_{i}\in U_{i}$, $v_{j}\in V_{j}$, and $w_{j}\in W_{j}$. For each $i$, let $C_{i}$ be a code of length $\ell$ over $F_{i}$, and for each $j$, let $C_{j}^{\prime}$ be a code of length $\ell$ over $G_{j}^{\prime}$ and let $C_{j}^{\prime\prime}$ be a code of length $\ell$ over $G_{j}^{\prime\prime}$. For $\mathbf{x_{i}}\in C_{i}$, $\mathbf{y_{j}^{\prime}}\in C_{j}^{\prime}$, and $\mathbf{y_{j}^{\prime\prime}}\in C_{j}^{\prime\prime}$, and for each $0\leq g\leq m-1$, let
@@
$2.$ The Duals of Quasi-Cyclic Codes
Recently, QC codes have been used in constructions of QECCs where orthogonality conditions on QC codes play a central role [Abdukhalikov-Bag-Panario2023, GalindoHernando, benjwal] . To study the duality of codes in $\mathbb{F}_{q}^{m\ell}$, we define the following inner products: Let $\textbf{a},\textbf{b}\in\mathbb{F}_{q}^{m\ell}$ where
@@
[lingsole, Proposition 3.2] Let $\mathbf{a},\mathbf{b}\in\mathbb{F}_{q}^{m\ell}$. Then $(T^{k\ell}(\mathbf{a}))\cdot\mathbf{b}=0$ for all $0\leq k\leq m-1$ if and only if $\langle\phi(\mathbf{a}),\phi(\mathbf{b})\rangle_{H}=0$, where $\phi$ is the aforementioned $R$-module isomorphism.
@@
[lingsole, Corollary 3.3] Let $\mathcal{C}$ be a quasi-cyclic code over $\mathbb{F}_{q}$ of length $m\ell$ and of index $\ell$ and let $\phi(\mathcal{C})$ be its image in $R_{q,m}^{\ell}$ under $\phi$. Then $\phi(\mathcal{C})^{\perp}=\phi(\mathcal{C}^{\perp})$, where the dual in $\mathbb{F}_{q}^{m\ell}$ is taken with respect to the Euclidean inner product, while the dual in $R_{q,m}^{\ell}$ is taken with respect to the Hermitian inner product. In particular, a quasi-cyclic code $\mathcal{C}$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if only if $\phi(\mathcal{C})$ is self-dual over $R_{q,m}$ with respect to the Hermitian inner product.
@@
[lingsole, Theorem 4.2] An $\ell$ quasi-cyclic code $\mathcal{C}$ of length $m\ell$ over $\mathbb{F}_{q}$ is self-dual with respect to the Euclidean inner product if and only if
@@
    As observed in [HKS, page 136], the irreducible factors $f_{i}$ which are self-reciprocal imply that the cardinality $q_{i}$ of each field extension $F_{i}$ is an even power of $q$, for all ii. However, there are two exceptions. The first one, for all $m$ and $q$, is the field extension coming from the irreducible factor x1x-1 of $x^{m}-1$. The other exception occurs when $q$ is odd and $m$ is even in which $x+1$ is another self-reciprocal factor of $x^{m}-1$. In both exceptions, $q_{i}=q$ and we equip these fields with the Euclidean inner product. Otherwise, we equip $F_{i}$ with the Hermitian inner product.
@@
The next result extends [lingsole, Theorem 4.2] to the dual of a QC code.
@@
[HKS, Proposition 7.3.5] Let $\mathcal{C}$ be a QC code with CRT decomposition as (10). Then its dual code $\mathcal{C}^{\perp}$ (under the Euclidean inner product) is of the form
@@
[HillNewton, Lemma 2.10] Suppose $C_{1}$ and $C_{2}$ are $n_{1},k,d_{1}]$ and $n_{2},k,d_{2}]$ codes over $\mathbb{F}_{q}$, respectively. If $C_{1}$ and $C_{2}$ have generator matrices $G_{1}$ and $G_{2}$, respectively, then $G_{1}|G_{2}]$ generates an $n_{1}+n_{2},k,d_{1}+d_{2}]$ code over $\mathbb{F}_{q}$ .
@@
In the special case when $C_{1}=C_{2}$, the code with generator matrix $G_{1}|G_{1}]$ will be denoted as $2C_{1}$, and such an argument may be naturally extended to $mC_{1}$, for any positive integer $m$. In this work, for any code $C$, we will refer to the code mCmC as $m$-copies of $C$ which will play an important role in the next section.